Grade 11Physical Chemistry

44 gram of a sample on complete combustion gives 88 gram CO2 and 36 gram of H2O. the molecular mass of the compound may be:A.C4H6B.C2H6OC.C2H4O

Samarth Raval
9 Years agoGrade 11

2 Answers

Vikas TU
9 Years ago
POAC on Carbon atom, for carbon compound CxHy,
x * 44/(12x + y) = 1*88/44
44x/(12x + y) = 2
22x = 12x + y
10x = y............(1)
POAC for H atom,
y*(44/(12x + y)) = 2*36/18
 
11y = 12x + y
10y = 12x
5y = 6x..........(2) 
Thus the formuale could be => CxH10x
Therefore it is possible in A option.
Prajwal
7 Years ago
88g CO2 means (88/44), 2moles of CO2
And 36g H2O means (36/18), 2moles of H2O
Thus, no. of moles of: C = 2, and H =2×2=4
Thus C and H in the molecular formula of compound is in ratio 1:2
The only option c has C and H in ratio 1:2
Hence correct answer is c